OpenStudy (anonymous):

If a leaky faucet wastes 45 gallons in three weeks, about how many gallons does it waste over a year (52 weeks)?

OpenStudy (anonymous):

52 / 3 = ?

hero (hero):

If it wastes 45 gallons in 3 weeks, It wastes 15 gallons per week on average. 15*52 = ?

OpenStudy (anonymous):

52/3 = ? . Take the answer you get for that and multiply it by 45 to get your answer

OpenStudy (anonymous):

2655 gallons :)

OpenStudy (anonymous):

I got a different answer @Aimee98

OpenStudy (anonymous):

What did you get then?

hero (hero):

http://lmgtfy.com/?q=15*52

OpenStudy (anonymous):

XD lol @Hero

OpenStudy (anonymous):

I got 780

OpenStudy (goformit100):

The Warmest Welcome to Open Study.`